find the material removal rate and time to machine.

To produce a hole with D=8mm and H=10mm at one end Al alloy rod by drilling. feed motion f=0.2mm/rev, rotational speed N is 1200rpm.

feed rate fr =N*F

Time to machine=(h+d/2)/fr

Material removal rate=πD^(2)fr/4
Answer #1

machining time = Le(equvalent lenght)/Vf(feed velocity)

Le= throught length + compulsory approach

Le = H + 0.5D(for through hole)

Le = 10+ 0.5*8 =14mm

Vf =fN =0.2*1200= 600mm/min

Machining time = 14/600 =0.0233 min =1.4s........

Material removal rate MRR= (π/4)D2*(fN)

MRR=(3.14/4)*82*(0.2*1200)=30144mm3/min=502.4mm3/s